# When Simultaneous Localization and Mapping Meets Wireless Communications: A Survey

### v1

@@ II Historical Background
The intersection of SLAM with wireless communications represents the convergence of two engineering domains that have evolved independently yet with complementary objectives. Although SLAM originates from robotics and autonomous navigation, wireless communications have long provided the infrastructure for connectivity and RF-based sensing. Examining their historical trajectories in parallel highlights how advances in navigation, signal processing, and communication progressively enabled more integrated positioning and mapping capabilities.
@@ III Foundations of SLAM in Networked Autonomous Systems
As autonomous systems expand beyond secluded operation, the assumptions that once defined SLAM as a self-contained estimation problem begin to break down. The environments in which these systems operate are increasingly structured by communication, computation, and control processes that interact continuously. In this setting, perception and estimation are no longer bounded by what a single platform can sense but are shaped by the information that flows through its internal and external networks. This shift transforms SLAM from a localized mapping task into a core mechanism that unifies perception and wireless communications with motion control, enabling a system to interpret and coordinate its behavior through shared and adaptive information. Therefore, this section investigates the underlying principles of a SLAM system and their interactions, linking measurement data acquisition and motion control processes. Specifically, it covers: Computer Vision Features in SLAM, which enable the extraction of landmarks and the surrounding scene structure, and provide robust priors that support accurate ego-motion estimation, Control in SLAM, being the fundamental mechanism of trajectory tracking, Passive SLAM, which derives spatial knowledge from signals and interactions that already permeate the environment, and Active SLAM, in which motion and coordination are deliberately guided by the SLAM system itself to improve landmark estimation and achieve the respective computer vision-driven goals such as detection confidence related and/or robotic surveillance-driven or communications-aware perception goals.
@@
A networked autonomous system can be deemed as a mobile, autonomous platform/vehicle, where observation, action, and communication form a unified process of continual feedback that dynamically shapes perception, reasoning, and interaction with the environment. As sensory information flows, state estimates and control signals strengthen each process in turn, forming adaptive cycles that bind perception, motion, and reasoning into a single dynamic loop [147]. Through these ongoing exchanges, the circulation of information keeps instantaneous control actions aligned with the system’s cognition, ensuring that every adjustment supports the stability of the overall state. Hence, autonomy arises from the sustained coordination that enables the system to respond coherently to both internal perturbations and external disturbances [101]. Formally, it can be defined by the following layers:
1. Interaction Layer: The interaction layer establishes the physical foundation of a networked autonomous system and forms the point where computational decisions take effect in the real world through motion. In this layer, control commands are converted into the forces and torques that generate and regulate movement while compensating for nonlinear behaviors and countering external influences such as surface roughness, longitudinal and lateral friction forces, aerodynamic drag, gravity, and hydrological dynamics. Through these regulating actions, the system maintains stable positioning with its surroundings and achieves the ability to move, manipulate, and interact in a controlled and adaptive manner.
2. Sensing Layer: The sensing layer captures how the system and its environment influence one another and transforms these interactions into measurable signals. Through optical, acoustic, radio and ranging modalities, the system observes and reconstructs external structures while compensating for variations in lighting, occlusion, or environmental dynamics. In connected settings, these sensing pathways also integrate information received through the communication stack, allowing perception to extend beyond direct observation and enhancing awareness through shared sensory data [24]. At the same time, internal motion is tracked through sensors that measure acceleration and body angular velocities. Together, these processes generate a continuous stream of data that describes both the vehicular dynamics and the optical/electromagnetic/mechanical responses of the surrounding structure. Thus, the Sensing Layer provides the perceptual channel through which computational intelligence begins to interpret the environment [117].
3. Estimation Layer: Based on sensory input, the estimation layer develops a coherent representation of the state of the system and the geometrical structure of its environment. Through the fusion of diverse measurement modalities, it retrieves vehicle kinematics whilst revealing the spatial organization of nearby primitives such as buildings, terrain, and communication infrastructure. SLAM functions as the central estimation framework, transforming data streams into evolving maps and trajectories that preserve consistency up to the retrievable scale over time, while the continual exchange of data through wireless links refines reconstructions and improves accuracy and stability in interconnected conditions [58]. The Estimation Layer thus converts observation into higher-level knowledge, establishing the basis from which reasoning and control emerge.
4. Inference Layer: Once a spatial model is formed, the inference layer interprets this knowledge to plan and execute purposeful actions. It connects perception to motion through decision-making, prediction, and high-level control processes that guide the system toward its objectives while adapting to environmental variation. Planning algorithms and predictive control frameworks within this layer generate trajectories, evaluate outcomes, and continuously adjust execution to maintain safety, efficiency, and stability. By translating spatial understanding into behavior, the Inference Layer closes the perception–action loop and prepares the system for coordinated operation within larger-scale connected contexts [147].
5. Coordination Layer: The coordination layer sustains coherence within a networked autonomous system by ensuring that perception, estimation, and control remain aligned through continuous communication across multiple vehicles/robotic platforms [31]. It governs how information, intent, and spatial knowledge are exchanged so that all subsystems share a consistent understanding of their operational environment. In practice, this coordination relies on low-latency and high-reliability communication frameworks that enable multi-agent estimation and control processes to remain synchronized even under dynamic conditions, as demonstrated in emerging 5G-enabled localization and cooperative perception systems [24, 150]. Through these mechanisms, local estimates converge to global objectives, supporting joint mapping, adaptive planning, and distributed reasoning [101]. The Coordination Layer thus closes the network-level feedback loop, embedding communication as a structural component of autonomy and binding perception, decision, and action into a single coherent process.
@@ III-A1 Computer Vision-aided SLAM General Outline
Classifying SLAM variants based on the nature of the sensing layer data outflow that is built upon, as well as the artificial intelligence primitives applied to them to extract patters, the first class that is prominent due to its great popularity in the Robotics community is V-SLAM. V-SLAM has emerged as a crucial technology for real-time localization and mapping, particularly in environments where external positioning infrastructure is not available. Unlike traditional SLAM methods that rely on LiDAR or inertia measurement units (IMUs), V-SLAM leverages camera-based sensing to extract features from the environment, match them across frames, and estimate motion through geometric and probabilistic models [143]. Key V-SLAM algorithms such as ORB-SLAM and LSD-SLAM have demonstrated robust performance across various scenarios, yet challenges persist, particularly in dynamic and unstructured environments [90]. V-SLAM algorithms generally consist of four main components: feature extraction, motion estimation, loop closure detection, and map optimization. Feature extraction is crucial for identifying and tracking distinctive points in the environment, while motion estimation determines the pose of the camera over time. Loop closure techniques improve accuracy by recognizing previously visited locations, reducing accumulated drift, and optimizing the overall map [90]. However, challenges such as occlusion, lighting variations, and real-time computational demands require continuous refinement of these components.
@@ III-B2 Low-Level Control and Estimation
Concerning the low-level control of autonomous vehicles, classic control techniques such as proportional-integral-derivative (PID) are commonly used. Apart from simple controllers, researchers have developed more sophisticated laws such as nonlinear control, adaptive control, sliding mode control, and neural network control to regulate a multitude of the vehicle’s degrees of freedom. Fuzzy logic controllers are popular control techniques. In [121], the sliding mode controller (SMC) and the fuzzy sliding mode controller (FSMC) were proposed to achieve the straight and turning motion of the robot. The feasibility and quality of the controllers were verified using numerical simulations. Aerial robots, when coupled with an IMU, can be controlled in an underactuated manner to perform agile flying maneuvers. The original quadrotor literature mostly focused on the modeling, state estimation, and low-level control of the vehicle while assuming that the localization, position, and orientation estimation required for feedback control is externally and readily available (e.g., from precise GNSS receivers or indoor motion capture systems). In fact, state estimators were eventually developed to close the control feedback loop without the need for external positioning-based devices, but this development occurred well after the control pipeline had been established [22]. The co-design occurs where the control influences the belief prediction step of a Gaussian State Estimator such as an EKF, often being central to the system that addresses SLAM as previously discussed.
@@ III-C Passive SLAM for Networked Autonomous Systems
Passive SLAM denotes the intrinsic ability of a networked autonomous system to infer spatial structure from the same data streams that sustain its operation, using wireless signals, feedback data, and shared measurements without deliberate exploration or additional probing transmissions [142]. System level-wide, this capability is realized through the combined operation of the sensing, state observation, and high-level control and coordination modules, where sensing captures environmental observations through visual and radio modalities, estimation fuses these into probabilistic spatial models that evolve with each update [90, 55], and the high-level control and coordination modules plan the path and motion trajectories whilst exchanging state information among distributed inter-vehicular or intra-vehicular modules [19]. Consequently, these coupled processes allow the networked autonomous system to maintain an accurate state estimate that remains stable under suitable modeling conditions even without explicit exploratory actions. Therefore, passive SLAM represents a fundamental attribute of networked autonomy, showing how continuous estimation and communication preserve collective spatial awareness throughout the system [101].
@@
During measurement data retrieval step of an autonomous system, wireless quantities such as time of arrival, angle of arrival, carrier phase, and multi-path structure, together with computer vision-related quaternions such as range, depth, and pose are exposed. Because these quantities characterize the relative geometry between the communicating agents and their environment, they can be directly incorporated into the estimation process without additional sensing effort. Hence, communication feedback signals that preserve connectivity also convey geometric information that supports spatial inference. This dual role is analogous to homography decomposition-based estimation in computer vision, where planar structures, inferred from dual-view feature correspondences related via a homography matrix, can be used to determine camera translational and angular velocities. Moreover, with some additional considerations for scale ambiguity resolution, the depth of the planar structure as well as a stabilizing velocity control law in the sense of Lyapunov Stability can be readily retrieved to align the structure with a desired reference frame[79]. Furthermore, the wireless propagation channel participates in estimation, since variations in signal properties reveal environmental geometry while reflecting network conditions such as latency and link reliability. Experimental and analytical studies in multipath-based SLAM confirm that reflections and virtual anchors contribute to joint trajectory and feature estimation, demonstrating that the propagation channel retains persistent geometric information about the scene [59]. In addition, research in decentralized SLAM shows that local maps can remain consistent when only selected loop closures or reduced factor-graph data are exchanged, with global alignment achieved through distributed optimization rather than centralized fusion [57]. Finally, developments in ISAC reveal that modern wireless systems are being designed to expose localization and channel-state information as intrinsic capabilities, embedding spatial inference directly within the operation of the network [72, 145]. Through these converging mechanisms, Passive SLAM emerges not as a separate sensing modality but as an embedded estimation process inherently sustained by the same feedback that enables control, coordination, and communication in a networked autonomous system.
@@
Attributing the aspects of perception-aware communication and communication-efficient mapping and localization methods, it has been demonstrated that spatial consistency can be preserved by transmitting only selected keyframes, compressed features, or reduced factor-graph structures, since each exchanged message is evaluated according to its information contribution rather than sent at a fixed rate [33]. Furthermore, adaptive estimation strategies show that update frequency, map resolution, and feature density can be adjusted in response to the instantaneous state of the wireless infrastructure and the available energy budget, allowing the networked autonomous system to maintain estimation stability even under fluctuating link quality [111].

### v2

@@ II Historical Background
The intersection of SLAM with wireless communications represents the convergence of two engineering domains that have evolved independently yet with complementary objectives. Although SLAM originates from robotics and autonomous navigation, wireless communications have long provided the infrastructure for connectivity and RF-based sensing. Examining their historical trajectories in parallel highlights how advances in navigation, signal processing, and communication progressively enabled more integrated positioning and mapping capabilities. Fig 1 shows a historical overview of significant developments in the intersection of SLAM and wireless communication technologies.
Figure 1: Timeline of significant developments in the intersection of SLAM and wireless communication technologies. Green indicates SLAM developments and blue the wireless communication technologies.
II-A Wireless Communications and localization evolution
@@ III Foundations of SLAM in Networked Autonomous Systems
As autonomous systems expand beyond secluded operation, the assumptions that once defined SLAM as a self-contained estimation problem begin to break down. The environments in which these systems operate are increasingly structured by communication, computation, and control processes that interact continuously. In this setting, perception and estimation are no longer bounded by what a single platform can sense but are shaped by the information that flows through its internal and external networks. This shift transforms SLAM from a localized mapping task into a core mechanism that unifies perception and wireless communications with motion control, enabling a system to interpret and coordinate its behavior through shared and adaptive information. Fig 2 illustrates a multilayered structure highlighting the linkages between components of a SLAM pipeline and related communications functions. Therefore, this section investigates the underlying principles of a SLAM system and their interactions, linking measurement data acquisition and motion control processes. Specifically, it covers: Computer Vision Features in SLAM, which enable the extraction of landmarks and the surrounding scene structure, and provide robust priors that support accurate ego-motion estimation, Control in SLAM, being the fundamental mechanism of trajectory tracking, Passive SLAM, which derives spatial knowledge from signals and interactions that already permeate the environment, and Active SLAM, in which motion and coordination are deliberately guided by the SLAM system itself to improve landmark estimation and achieve the respective computer vision-driven goals such as detection confidence related and/or robotic surveillance-driven or communications-aware perception goals.
Figure 2: A multilayered, sensor-agnostic structure highlighting the linkages between components of a SLAM pipeline and communications functions where bidirectional relationships are noted as well as the convergence in communications-driven active perception is outlined.
@@
A networked autonomous system can be deemed as a mobile, autonomous platform/vehicle, where observation, action, and communication form a unified process of continual feedback that dynamically shapes perception, reasoning, and interaction with the environment. As sensory information flows, state estimates and control signals strengthen each process in turn, forming adaptive cycles that bind perception, motion, and reasoning into a single dynamic loop [150]. Through these ongoing exchanges, the circulation of information keeps instantaneous control actions aligned with the system’s cognition, ensuring that every adjustment supports the stability of the overall state. Hence, autonomy arises from the sustained coordination that enables the system to respond coherently to both internal perturbations and external disturbances [103]. Formally, it can be defined by the following layers that are also illustrated in Fig. 3:
@@
1. Interaction Layer: The interaction layer establishes the physical foundation of a networked autonomous system and forms the point where computational decisions take effect in the real world through motion. In this layer, control commands are converted into the forces and torques that generate and regulate movement while compensating for nonlinear behaviors and countering external influences such as surface roughness, longitudinal and lateral friction forces, aerodynamic drag, gravity, and hydrological dynamics. These functions can be realized through low-level control techniques such as proportional–integral–derivative (PID) control, model predictive control, and nonlinear feedback linearization control, enabling efficient trajectory tracking under time-varying dynamics. Through these regulating actions, the system maintains stable positioning with its surroundings and achieves the ability to move, manipulate, and interact in a controlled and adaptive manner.
2. Sensing Layer: The sensing layer captures how the system and its environment influence one another and transforms these interactions into measurable signals. Through optical, acoustic, radio and ranging modalities, the system observes and reconstructs external structures while facing variations in lighting, occlusion, or environmental dynamics. Key functions may include ranging, extraction of RF/optical features, extraction of semantic context, and determination of acceleration/angular velocity, which are enabled through techniques such as fast Fourier transform (FFT)-based radar signal processing, camera-based feature detection, LiDAR point cloud processing, and inertial sensing. In connected settings, these sensing pathways also integrate information received through the communication stack, allowing perception to extend beyond direct observation and enhancing awareness through shared sensory data [25]. Together, these processes generate a continuous stream of data that describes both the vehicular dynamics and the optical/electromagnetic/mechanical responses of the surrounding structure. Thus, the Sensing Layer provides the perceptual channel through which computational intelligence begins to interpret the environment [119].
3. Estimation Layer: Based on sensory input, the estimation layer develops a coherent representation of the state of the system and the geometrical structure of its environment. Through the fusion of diverse measurement modalities, it retrieves vehicle kinematics whilst revealing the spatial organization of nearby primitives such as buildings, terrain, and communication infrastructure. In this layer, state estimation uncertainty quantification and map construction can be approached through probabilistic techniques such as Bayesian filtering (e.g., EKF, particle filter), iterative optimization, and factor graph methods. SLAM functions as the central estimation framework, transforming data streams into evolving maps and trajectories that preserve consistency up to the retrievable scale over time, while the continual exchange of data through wireless links refines reconstructions and improves accuracy and stability in interconnected conditions [60]. The Estimation Layer thus converts observation into higher-level knowledge, establishing the basis from which reasoning and control emerge.
4. Inference Layer: Once a spatial model is formed, the inference layer interprets this knowledge to plan and execute purposeful actions. It connects perception to motion through decision-making, prediction, and high-level control processes that guide the system toward its objectives while adapting to environmental variations. Its key capabilities can include trajectory planning, risk-aware decision-making, and high-level control, typically achieved through optimization-based motion planning algorithms, sampling-based motion planning algorithms such as the rapidly exploring random tree (RRT), and deep reinforcement learning-based policies. Planning algorithms and predictive control frameworks within this layer generate trajectories, evaluate outcomes, and continuously adjust execution to maintain safety, efficiency, and stability. By translating spatial understanding into behavior, the Inference Layer closes the perception–action loop and prepares the system for coordinated operation within larger-scale connected contexts [150].
5. Coordination Layer: The coordination layer sustains coherence within a networked autonomous system by ensuring that perception, estimation, and control remain aligned through continuous communication across multiple vehicles/robotic platforms [32]. It governs how information, intent, and spatial knowledge are exchanged so that all subsystems share a consistent understanding of their operational environment. Key functions may include distributed state estimation and networked control under communication constraints, which can be enabled through techniques such as cooperative perception and localization, and event-triggered communication strategies. In practice, this coordination relies on low-latency and high-reliability communication frameworks that enable multi-agent estimation and control processes to remain synchronized even under dynamic conditions, as demonstrated in emerging 5G-enabled localization and cooperative perception systems [25, 153]. From an information-theoretic perspective, the coordination layer is linked to the relationship between communication rate and system stability, where sufficient information flow needs to be maintained to counteract unstable modes in the system dynamics, as formalized by the data-rate theorem in networked control systems [32]. Through these mechanisms, local estimates and control actions converge to global objectives, supporting joint mapping, adaptive planning, and distributed reasoning [103]. The Coordination Layer thus closes the network-level feedback loop, embedding communication as a structural component of autonomy and binding perception, decision, and action into a single coherent process.
@@ III-A1 Computer Vision-aided SLAM General Outline
Classifying SLAM variants based on the nature of the sensing layer data outflow that is built upon, as well as the artificial intelligence primitives applied to them to extract patters, the first class that is prominent due to its great popularity in the Robotics community is V-SLAM. V-SLAM has emerged as a crucial technology for real-time localization and mapping, particularly in environments where external positioning infrastructure is not available. Unlike traditional SLAM methods that rely on LiDAR or inertia measurement units (IMUs), V-SLAM leverages camera-based sensing to extract features from the environment, match them across frames, and estimate motion through geometric and probabilistic models [145]. Key V-SLAM algorithms such as ORB-SLAM and LSD-SLAM have demonstrated robust performance across various scenarios, yet challenges persist, particularly in dynamic and unstructured environments [93]. V-SLAM algorithms generally consist of four main components: feature extraction, motion estimation, loop closure detection, and map optimization. Fig. 4 shows the components that can be found in a V-SLAM pipeline [36]. Feature extraction is crucial for identifying and tracking distinctive points in the environment, while motion estimation determines the pose of the camera over time. Loop closure techniques improve accuracy by recognizing previously visited locations, reducing accumulated drift, and optimizing the overall map [93]. However, challenges such as occlusion, lighting variations, and real-time computational demands require continuous refinement of these components.
Figure 4: Illustration of the front-end and back-end functionalities in a V-SLAM pipeline. Image frames and Depth Maps are based on the KITTI dataset [36]
@@ III-B2 Low-Level Control and Estimation
Concerning the low-level control of autonomous vehicles, classic control techniques such as PID control are commonly used. Apart from simple controllers, researchers have developed more sophisticated laws such as nonlinear control, adaptive control, sliding mode control, and neural network control to regulate a multitude of the vehicle’s degrees of freedom. Fuzzy logic controllers are popular control techniques. In [123], the sliding mode controller (SMC) and the fuzzy sliding mode controller (FSMC) were proposed to achieve the straight and turning motion of the robot. The feasibility and quality of the controllers were verified using numerical simulations. Aerial robots, when coupled with an IMU, can be controlled in an underactuated manner to perform agile flying maneuvers. The original quadrotor literature mostly focused on the modeling, state estimation, and low-level control of the vehicle while assuming that the localization, position, and orientation estimation required for feedback control is externally and readily available (e.g., from precise GNSS receivers or indoor motion capture systems). In fact, state estimators were eventually developed to close the control feedback loop without the need for external positioning-based devices, but this development occurred well after the control pipeline had been established [23]. The co-design occurs where the control influences the belief prediction step of a Gaussian State Estimator such as an EKF, often being central to the system that addresses SLAM as previously discussed.
@@ III-C Passive SLAM for Networked Autonomous Systems
Passive SLAM denotes the intrinsic ability of a networked autonomous system to infer motion and spatial structure from the same data streams that sustain its operation, using vision data, wireless signals, and shared measurements without deliberate exploration or additional probing transmissions [144]. System level-wide, this capability is realized through the combined operation of the sensing, state observation, and high-level control and coordination modules, where sensing captures environmental observations through visual and radio modalities, estimation fuses these into probabilistic spatial models that evolve with each update [93, 57], and the high-level control and coordination modules plan the path and motion trajectories whilst exchanging state information among distributed inter-vehicular or intra-vehicular modules [21]. Consequently, these coupled processes allow the networked autonomous system to maintain an accurate state estimate that remains stable under suitable modeling conditions even without explicit exploratory actions. Therefore, passive SLAM represents a fundamental attribute of networked autonomy, showing how continuous estimation and communication preserve collective spatial awareness throughout the system [103].
@@
During measurement data acquisition step of an autonomous system, wireless quantities such as time of arrival, angle of arrival, carrier phase, and multi-path structure, together with computer vision-related quantities such as range, depth, and pose are retrieved. Because these quantities characterize the relative geometry between the agents and their environment, they can be directly incorporated into the estimation process without additional sensing effort. Hence, communication feedback signals that preserve connectivity also convey geometric information that supports spatial inference. An interesting capability they may encompass is providing the source of metric information that aids scale recovery in the context of monocular V-SLAM. In a general setting, their dual role of carrying information and embedding geometry is analogous to the estimation based on essential matrix decomposition in computer vision, where 3-D landmarks, inferred from multi-view feature correspondences related via the epipolar geometry, can be used to determine camera translational and angular velocities[108]. Furthermore, the wireless propagation channel participates in estimation, since variations in signal properties reveal environmental geometry while reflecting network conditions such as latency and link reliability. Experimental and analytical studies in multipath-based SLAM confirm that reflections and virtual anchors contribute to joint trajectory and feature estimation, demonstrating that the propagation channel retains persistent geometric information about the scene [61]. In addition, research in decentralized SLAM shows that local maps can remain consistent when only selected loop closures or reduced factor-graph data are exchanged, with global alignment achieved through distributed optimization rather than centralized fusion [59]. Finally, developments in ISAC reveal that modern wireless systems are being designed to expose localization and channel-state information as intrinsic capabilities, embedding spatial inference directly within the operation of the network [74, 148].
@@
Attributing the aspects of perception-aware communication and communication-efficient mapping and localization methods, it has been demonstrated that spatial consistency can be preserved by transmitting only selected keyframes, compressed features, or reduced factor-graph structures, since each exchanged message is evaluated according to its information contribution rather than sent at a fixed rate [34]. Furthermore, adaptive estimation strategies show that update frequency, map resolution, and feature density can be adjusted in response to the instantaneous measurements of the wireless infrastructure and the available energy budget, allowing the networked autonomous system to maintain estimation stability even under fluctuating link quality [113].
To provide a unified mathematical framework where pose estimation, motion input, and geometric channel-related quantities are integrated into a single estimation model, the problem of V-SLAM for a vehicle equipped with a monocular camera operating in an urban environment is considered. Let $I_{k-1}$ and $I_{k}$ denote two consecutive image frames acquired at discrete time instants k1k-1 and $k$.
III-C1 Feature Detection and Tracking
At initialization ($k=0$), the pixel coordinates of salient features in the image signal that can be extracted using a geometric primitive detection method, denoted (per-feature) by
$\mathbf{x}_{i,0}=\begin{bmatrix}x_{i,0},\ y_{i,0}\end{bmatrix}^{\top}$. (1)
Feature detection may occur at the beginning of the autonomous vehicle navigation process. The future pixel location of each feature in the subsequent image can be predicted using a feature tracking algorithm. The Lucas-Kanade (LK) optical flow is a tracking algorithm known for its efficiency realizing the pyramid sampling mechanism for coarse-to-fine estimation as well as its real-time performance [81, 15]. According to this algorithm, for each feature $i$, its displacement $\mathbf{d}_{i,k}$ is obtained by minimizing:
$\mathbf{d}_{i,k}=\arg\min_{\mathbf{d}}\sum_{\mathbf{x}\in M_{i}}\left(I_{k}(\mathbf{x}+\mathbf{d})-I_{k-1}(\mathbf{x})\right)^{2},$ (2)
where $M_{i}$ denotes a local image patch. It is remarked that constant brightness is assumed between two consecutive frames. The tracked feature location is then
$\mathbf{x}_{i,k}=\mathbf{x}_{i,k-1}+\mathbf{d}_{i,k}.$ (3)
The initialization of the features to track with the LK algorithm can also occur at preselected fixed time intervals or upon loss of the tracks.
III-C2 Epipolar Geometry and Relative Pose Estimation
Feature points corresponding to the same 3-D structure, as tracked from two consecutive camera viewpoints are constrained each to be in the epipolar plane formed by the two locations of the camera center and the respective 3-D point. The latter intersects each image plane at the two measurement acquisition instances. Solving the epipolar constraint can lead to the retrieval of the pose of the autonomous vehicle, under certain assumptions.
To derive the epipolar constraint model, the image coordinates of a tracked feature at time $k$ are normalized by the horizontal and vertical focal lengths $f_{x}$ and $f_{y}$, respectively, yielding:
$\tilde{\mathbf{x}}_{i,k}=\begin{bmatrix}(x_{i,k}-c_{x})/f_{x},\ (y_{i,k}-c_{y})/f_{y},\ 1\end{bmatrix}^{\top}$. (4)
$c_{x}$ and $c_{y}$ correspond to the pixel coordinates of the principal point, often taken as the center of the image.
The essential matrix $\mathbf{E}_{k}$ satisfies the epipolar constraint:
$\tilde{\mathbf{x}}_{i,k}^{\top}\mathbf{E}_{k}\tilde{\mathbf{x}}_{i,k-1}=0.$ (5)
Using robust parameter estimation methods and singular value decomposition (SVD), $\mathbf{E}_{k}$ is decomposed as:
$\mathbf{E}_{k}=[\hat{\mathbf{t}}_{k}]_{\times}\hat{\mathbf{R}}_{k},$ (6)
yielding the relative rotation in the special orthogonal group $(3)$ $\hat{\mathbf{R}}_{k}\in SO(3)$ and the translation direction $\hat{\mathbf{t}}_{k}\in\mathbb{R}^{3}$, defined up to an unknown scale. For brevity, $\hat{\mathbf{R}}_{k}$ and $\hat{\mathbf{t}}_{k}$ denote the relative rotation and translation from the camera coordinate frame at k1k-1 to the frame at $k$.
The corresponding homogeneous transformation is:
$\hat{\mathbf{H}}_{k}=\begin{bmatrix}\hat{\mathbf{R}}_{k}&\hat{\mathbf{t}}_{k}\\ \mathbf{0}^{T}&1\end{bmatrix}$. (7)
Pose estimates w.r.t the initial camera coordinate system, considered the inertial frame hereinafter, are obtained recursively:
$\mathbf{P}_{k}=\mathbf{P}_{k-1}\hat{\mathbf{H}}_{k}^{-1}.$ (8)
III-C3 Landmark Reconstruction
Given the feature correspondences across camera views, the apparent motion of a 3-D landmark $\mathbf{X}_{i}$, as viewed from a moving camera, is:
$\mathbf{\dot{X}}_{i}(t)=\mathbf{\Omega}(t)\mathbf{X}_{i}(t)+\mathbf{v}(t).$ (9)
The discrete-time form is:
$\mathbf{X}_{i,k}=e^{\mathbf{\Omega}_{k}dt}\mathbf{X}_{i,k-1}+\mathbf{v}_{k}dt,$ (10)
where $\mathbf{\Omega}_{k}=[\boldsymbol{\omega}_{k}]_{\times}$ is the skew-symmetric matrix in $SO(3)$ that describes the angular velocity the camera exerted, responsible for the change of orientation from the coordinate frame at k1k-1 to the coordinate frame at $k$, and expressed in the former coordinate frame. $\mathbf{v}_{k}$ represents the translational velocity of the camera from the coordinate frame at k1k-1 to the coordinate frame at $k$, expressed in the former coordinate frame.
Using the rotation and translation components of the pose estimate and exploiting the fact that $\hat{\mathbf{R}}_{k}\approx e^{\mathbf{\Omega}_{k}dt}$ and $\hat{\mathbf{t}}_{k}\approx\frac{1}{\lambda_{k}}\mathbf{v}_{k}dt$:
$\hat{\mathbf{X}}_{i,k}=\hat{\mathbf{R}}_{k}\hat{\mathbf{X}}_{i,k-1}+\hat{\mathbf{t}}_{k},$ (11)
where $\hat{\mathbf{X}}_{i,k}=\frac{1}{\lambda_{k}}\mathbf{X}_{i,k}$ represents the location of each tracked landmark up to an arbitrary scale. Scale recovery followed by the determination of the landmarks’ locations w.r.t an inertial coordinate frame results in a metric raw map.
III-C4 Scale Ambiguity and Metric Motion Input
The essential matrix decomposition provides the relative translation up to scale:
$\mathbf{t}_{k}=\lambda_{k}\hat{\mathbf{t}}_{k},$ (12)
where $\lambda_{k}>0$ is an unknown factor.
To resolve this ambiguity, one can acquire via the odometry sensors the following metric displacement that, in the dynamics of both the autonomous vehicle and the landmarks as seen from the moving camera, serves the role of the motion control input [118]:
$\Delta\mathbf{p}_{k}\in\mathbb{R}^{3},$ (13)
providing the true translational motion between frames in metric units. $\lambda_{k}$ is estimated by solving:
$\lambda_{k}=\arg\min_{\lambda>0}\left\|\Delta\mathbf{p}_{k}-\lambda\hat{\mathbf{t}}_{k}\right\|_{2}^{2}.$ (14)
Once $\lambda_{k}$ is obtained, the pose and the landmarks can be expressed in metric units.
III-C5 Sources of Metric Motion Input
The metric displacement $\Delta\mathbf{p}_{k}$ can be obtained from different sensing modalities:
Inertial Odometry
Inertial measurements provide acceleration $\mathbf{a}_{k}$, from which velocity and displacement can be estimated via double integration, numerically approximated as:
$\Delta\mathbf{p}_{k}\approx\mathbf{v}_{k}\Delta t.$ (15)
$\mathbf{v}_{k}=\mathbf{v}_{k-1}+\mathbf{a}_{k}\Delta t,$ (16)
This procedure is often prone to drift, raising the need for sensor fusion (e.g. through Kalman filtering) along with bias compensation and gravity removal.
Wireless Communications-based Odometry
Metric position estimates can alternatively be obtained from communication signals exchanged with multiple base stations (gNBs) of known positions $\mathbf{s}_{b}$. Measurements such as time-of-arrival (ToA), round-trip time (RTT), angle-of-arrival (AoA), and received signal strength (RSS) can satisfy the following measurement model [48]:
$z_{b,k}=h_{b}(\mathbf{p}_{k},\mathbf{s}_{b})+n_{b,k},$ (17)
where $\mathbf{p}_{k}$ denotes the position of the autonomous vehicle and $n_{b,k}$ measurement noise.
By combining measurements from multiple gNBs, a position estimate $\mathbf{p}_{k}^{\mathrm{RF}}$ can be obtained via multi-lateration/angulation [48]. The corresponding metric displacement is:
$\Delta\mathbf{p}_{k}^{\mathrm{RF}}=\mathbf{p}_{k}^{\mathrm{RF}}-\mathbf{p}_{k-1}^{\mathrm{RF}}.$ (18)
This quantity could be used directly in the scale estimation problem:
$\lambda_{k}=\arg\min_{\lambda>0}\left\|\Delta\mathbf{p}_{k}^{\mathrm{RF}}-\lambda\hat{\mathbf{t}}_{k}\right\|_{2}^{2}.$ (19)